# In-medium bottomonium properties from lattice NRQCD calculations with extended meson operators

### v1

@@ III.1 Continuum-subtracted effective masses and in-medium bottomonium properties from simple ansätze
which for very large and very small $x$ behaves as $\exp(-2x/\delta)$. In the limit $\delta\rightarrow 0$, we recover the cut Lorentzian form. The parametrization of the spectral function given by Eq. 7, Eq. 8, and Eq. 10, which we will call it smooth-cut Lorentizan, can describe the spectral function obtained in the T-matrix approach Tang et al. (2024b). This is shown in Fig. 8 for a 1P bottomonium state as an example. The parameter $\Gamma_{0}$ was obtained as the width at half maximum of the T-matrix spectral function, and the parameters $cut$ and $\delta$ were adjusted to obtain the best possible description of the T-matrix spectral functions. We also find that the Lorentzian form can describe the T-matrix spectral function very well in the interval $[M-{cut},M+{cut}]$, where ${cut}\simeq 4\Gamma_{0}$. A similar statement holds true for the spectral function of the static quark-antiquark pair obtained in hard thermal loop perturbation theory Burnier and Rothkopf (2013).
@@
The in-medium mass shift $\Delta M(T)$ from wave-function-optimized operators is shown in Fig. 7. Interestingly, we see no mass shift within estimated errors for both the smooth-cut Lorentzian form and the cut Lorentzian form, for both lattice spacings. This is consistent with the previous findings in Refs. Larsen et al. (2019, 2020a) and also means that the absence of the mass shift is a robust result independent from the detailed form of the spectral spectral function. We elaborate on this in Appendix B. The $\Delta M(T)$ for $\Upsilon$(1S) and $\chi_{b0}$(1P) states obtained from the correlators using Gaussian sources agree with the ones shown in Fig. 7 within errors.
@@
The in-medium width increases with the temperature and is larger for higher-lying states, i.e., it follows a hierarchy pattern corresponding to the hierarchy of bottomonium masses and sizes. In the case of the $\Upsilon$(1S), $\Upsilon$(2S), and $\chi_{b0}$(1P) states, the estimates of the thermal width obtained using the cut Lorentzian are significantly larger than the ones obtained using the smooth-cut Lorentzian. The difference is the largest for the $\Upsilon$(1S) state, where it is as large as a factor three. The difference in $\Gamma_{\rm med}(T)$ obtained from the two forms of the spectral function implies that the tail structure of the quasi-particle peak plays an important role in the observed decrease in the continuum-subtracted effective mass. The smaller the width, the larger the contribution from the tails of the dominant peak to the decrease in the effective masses. For this reason, the difference in $\Gamma_{\rm med}$ from using the two ansätze is the largest for the $\Upsilon$(1S). For the $\Upsilon$(3S) and the $\chi_{b0}$(2P), the difference in thermal width arising from the use of the two ansätze is much smaller, and for the $\chi_{b0}$(2P) state it is almost statistically insignificant. The reason for this is the following. When the width is large, the decrease in the effective masses is largely determined by the behavior of the spectral function around the peak, as explained in the Appendix B, and the tails of the quasi-particle peak do not contribute much to the observed decrease of the continuum-subtracted effective masses. The thermal widths obtained in this work are significantly smaller than the ones obtained in Refs. Larsen et al. (2019, 2020a). The reason for this is the use of a simple Gaussian form for the quasi-particle peak in those studies, which is not physically motivated. For a Gaussian form of the peak, the slope of the continuum-subtracted effective masses is entirely determined by the width parameter. As discussed above, this is not the case for realistic spectral functions, for which the shape of the spectral function far away from the peak position can largely contribute to the slope of the continuum-subtracted effective masses.
@@ IV Conclusions
In this paper we studied the in-medium properties of bottomonium using NRQCD with extended meson operators at two lattice spacings, $a=0.0602$ fm and $a=0.0493$ fm. We used temporal lattice extents $N_{\tau}=30-16$, covering the temperature range $T=133-251$ MeV. By studying the temperature dependence of bottomonium correlation functions, as well as through the comparison of the finite-temperature correlation function to the zero-temperature correlation function, we proposed a physically motivated parametrization of the in-medium spectral function that contains a dominant quasi-particle peak corresponding to the bottomonium states of interest. We found that all bottomonium states below the open-beauty threshold can exist as well-defined quasi-particle excitations up to $T=251$ MeV, though some of these excitations have large thermal width. For temperature $T\lesssim 167$ MeV the existence of bottomonium states and the values of their masses can be established without relying on any specific form of the spectral function. For the $\Upsilon$(1S) this is true even for $T\lesssim 182$ MeV. To obtain the in-medium bottomonium mass and width at higher temperatures, we considered several parametrizations of the dominant quasi-particle peak. We found that, irrespective of the details of these parametrizations, the in-medium bottomonium masses agree with the corresponding vacuum masses within errors. This is in contrast to the usual picture of bottomonium in quark-gluon plasma based on color screening, where there is significant shift in bottomonium masses and higher excited states are melted at temperatures somewhat above the crossover temperature.
@@
The estimates of the thermal width are quite sensitive to the detailed spectral shape of the quasi-particle peak. We considered two physically motivated choices of this spectral shape to obtain the thermal widths, which are shown in Fig. 9. These estimates of the thermal width are considerably smaller than the previous estimates from lattice NRQCD calculations that used a Gaussian form of the dominant peak Larsen et al. (2020a, 2019), especially for the $\Upsilon$(1S). However, the two estimates of the thermal widths differ significantly, in particular for the $\Upsilon$(1S) state, where the differences can reach up to a factor of three. For the excited states, the differences in the thermal width obtained from the different spectral shapes are smaller, but, in general, estimating the thermal bottomonium width on the lattice better than a factor of two appears to be very challenging. Nevertheless, it would be very interesting to see how our estimates of the thermal width impact the phenomenological description of the nuclear suppression factor, $R_{AA}$, of the $\Upsilon$(1S), $\Upsilon$(2S), and $\Upsilon$(3S) states in heavy-ion collisions.
@@ Acknowledgements.
This work is supported partly by the National Key Research and Development Program of China under Contract No. 2022YFA1604900; the National Natural Science Foundation of China under Grants No. 12293064, No. 12293060 and No. 12325508, and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704, through Heavy Flavor Theory for QCD Matter (HEFTY) topical collaboration in Nuclear Theory, and within the framework of Scientific Discovery through Advance Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond. S. Meinel acknowledges support by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009913. W.-P. Huang acknowledges support from the Young Elite Scientists Sponsorship Program (Special Support for Doctoral Students) by CAST. This research used awards of computer time provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02- 05CH11231, on Frontier supercomputer in Oak Ridge Leadership Class Facility through ALCC Award, and the awards on JUWELS at GCS@FZJ, Germany and Marconi100 at CINECA, Italy. Computations for this work were carried out in part on facilities of the USQCD Collaboration, which are funded by the Office of Science of the U.S. Department of Energy. Some of the calculations were performed on the GPU cluster in the Nuclear Science Computing Center at CCNU (NSC3) and Wuhan supercomputing center.
@@ Appendix A DETAILS OF MASS EXTRACTIONS IN VACUUM
In this Appendix, we describe the detailed procedure of fitting the vacuum correlators with an exponential ansätze to extract the corresponding mass parameters. We fitted the correlators to a single-exponential form within a range where the ground-state contribution dominates. Thus, careful selection of the fit interval is essential to eliminate contamination from higher states while maintaining statistical precision. To achieve this, we followed a procedure adapted from Refs. Fritzsch et al. (2012); Bahr (2015). This procedure is exemplified in Fig. 10, using the $\chi_{b0}$(1P) case as an illustration. The upper end of the fit interval $\tau_{\text{max}}$ is determined based on the temporal separation where the signal-to-noise ratio of the effective mass $M_{\mathrm{eff}}(\tau)$ reaches a specified threshold, with the effective mass defined as $M_{\mathrm{eff}}(\tau)=-\partial_{\tau}\log C(\tau)$. In our analyses, we set a conservative criterion of $1\%$ for measurements employing Gaussian-smearing sources and $8\%$ for those using wave-function-optimized sources.
@@
To determine the lower end of the fit interval $\tau_{\text{min}}$, we initially performed a two-exponential fit to the correlators within a range where the two-state ansätze $f_{n}(\tau)=\sum_{i=0}^{n}A_{i}e^{-E_{i}\tau}$ with $n=1$ described the data well. This is illustrated by the solid black line in Fig. 10, with the two-state fit performed outside the gray-shaded area and ending at $\tau_{\text{max}}$. The fit parameters $A_{n=0,1}$ and $E_{n=0,1}$ were then used to reconstruct the excited-state contribution to the effective mass as $\delta M(\tau)=\log[f_{1}(\tau)/f_{1}(\tau+a)]-\log[f_{0}(\tau)/f_{0}(\tau+a)]$. We set $\tau_{\text{min}}$ to the smallest $\tau$ value where $\delta M(\tau)/E_{0}<25\%\times\delta_{M_{\mathrm{eff}}}(\tau)/M_{\mathrm{eff}}(\tau)$ is satisfied, with $\delta_{M_{\mathrm{eff}}}(\tau)$ representing the statistical uncertainty of $M_{\mathrm{eff}}$ at $\tau$. In this way, from $\tau_{\text{min}}$, the excited-state contamination (or, equivalently, the systematic uncertainty due to the fit ansätze) falls below the statistical uncertainty.
@@
For the $\chi_{b0}$(1P) case, $\tau_{\text{min}}$ is indicated by the dashed line in Fig. 10, with the interval $[\tau_{\text{min}},\tau_{\text{max}}]$ marked as the blue band. Within this interval, the ground state dominates, allowing for a single-exponential fit to extract the information of the ground state. The error of the extracted mass is estimated using the Bootstrap method. The final central value of the mass was the median of the Bootstrap ensemble of fit results, with the error quoted as the 1-$\sigma$ width of this distribution. It is worth mentioning that for quantities such as mass differences, which depend on results from different fits, each fit was performed on the same Bootstrap sample. The final quantity for one sample is then computed using these fit results from the same sample, giving an approximate probability distribution. The mean value and the 1-$\sigma$ width of this distribution are then quoted.
Figure 10: The effective mass for $\chi_{b0}$(1P) (red points) and the mass (blue band) extracted from a single-exponential fit to the corresponding correlator, both in lattice units. To determine the lower bound $\tau_{\text{min}}$ of the fit, initially, a two-state fit (black line) to the correlator is performed outside the gray-shaded area and ending at $\tau_{\text{max}}$. The determined interval $[\tau_{\text{min}},\tau_{\text{max}}]$ for the single-exponential fit is shown as the range of the blue band.
@@ Appendix B EXTRACTIONS OF IN-MEDIUM PARAMETERS AT FINITE TEMPERATURES
As discusssed in Sec. III of the main text, the in-medium bottomonium parameters such as the widths and masses are encoded in the continuum-subtracted in-medium spectral function $\rho_{\rm med}(\omega,T)$. Furthermore, $\rho_{\rm med}(\omega,T)$ includes a quasi-particle peak containing information on the in-medium bottomonium masses and widths, and also a small contribution at frequencies well below the dominant peak controlling the behavior of the correlator at large $\tau$ around $1/T$. For the first part, we use the physically motivated cut Lorentzian ansätze or the smooth-cut Lorentzian ansätze, as discussed in the main text, while for the other part at $\omega$ well below the dominant-peak position, a single delta function is sufficient to describe the correlators at large $\tau$, shown as the first term of Eq. 11. In the main text it was argued that, for the cut Lorentzian, a reasonable choice for the cut is $4\Gamma_{0}$. Then an important question to address is how the obtained width depends on the choice of the parameter $cut$. A smaller value of $cut$ will result in a larger width. On the other hand the value of $cut$ cannot be too small because that would mean that the quasi-particle peak is not well defined, as other parts the spectral function provide a much larger contribution to the correlator. The value ${cut}=2\Gamma_{0}$ is probably the smallest value compatible with the existence of a quasi-particle peak, as the spectral function around the quasi-particle peak will decrease by a large factor before other structure in the spectral function becomes dominant. Therefore, we also considered the choice ${cut}=2\Gamma_{0}$ in our analysis as discussed below. For sufficiently low temperature, the thermal width can be too small to be resolved, given our statistical errors. In such cases, it makes sense to set the thermal width to zero, i.e., to assume a delta function for the quasi-particle peak. For sufficiently low temperature, the small contribution below the quasi-particle peak also does not play a role and can be neglected. This corresponds to fitting the effective masses by a constant. The results of these fits are shown in Tab. 2 and Tab. 3 for the $\Upsilon$ and $\chi_{b0}$ states respectively. In the cases of the $\Upsilon$(1S) state, the constant fit has worse $\chi^{2}/d.o.f$ compared to other fits only for $T\gtrsim 182$ MeV, while for other bottomonium states the constant fit has worse $\chi^{2}/d.o.f$ only for $T\gtrsim 163$ MeV.
@@
We also performed fits using the Gaussian form of the quasi-particle peak in the spectral function given by Eq. 6 to make contact with previous analyses Larsen et al. (2020a, 2019), which are summarized in Tab. 6.
@@
The in-medium widths and mass shifts from different types of parametrizations of the in-medium spectral function $\rho_{\rm med}(\omega,T)$ are shown in Fig. 11 and Fig. 12. It can be seen from the comparison of open and filled points with the same shape in Fig. 11 that, for the excited bottomonium states, the widths extracted using the ansätze excluding a delta-function term tend to be systematically larger than those including the delta function, no matter what type of ansätze is used. This is especially obvious in the cases of high temperatures and for excited states. This indicates that the minor structure below the dominant peak of the spectral function matters for these cases, as expected. This delta peak is not well separated from the main peak structure of the spectral function, and interferences between them are possible. Thus, the tail of the main peak will compensate for the delta peak contribution if the delta term is excluded during the fits, leading to a larger width. In the case of the Gaussian fits the thermal width in this is defined as the width at the half maximum, which is related to the Gaussian width parameters as $\sqrt{2\ln 2}\Gamma_{G}$ Shi et al. (2022). With this definition the thermal width from Gaussian fits width agrees with the cut Lorentzian results with $cut=2\Gamma_{0}$. As the cut increases to $4\Gamma_{0}$, the width extracted from cut Lorentzian ansätze decreases to about 50% $\sim$ 60% of the $cut=2\Gamma_{0}$ result, also stressing the importance of modeling the tail of the main peak structure of the spectral function. We also compare our estimates for the thermal width obtained from the cut Lorentzian form to estimates using the Gaussian ansätze for the dominant peak obtained in this analysis as well as in Ref. Larsen et al. (2019). Here again the width at half maximum was used to define the thermal width. Since the choice $cut=2\Gamma_{0}$ is the smallest value of cut still compatible with a well-defined quasi-particle peak and leads to the larger thermal width, the estimate of the width obtained from the Gaussian form of the dominant peak can be considered as an upper bound on the thermal bottomonium width. The in-medium mass shifts in Fig. 12 are all consistent with zero for all states under our consideration, irrelevant of the parametrizations of $\rho_{\rm med}(\omega,T)$.
@@
The detailed choice of the quasi-particle peak has significant effect on the estimated thermal bottomonium width, while the position of the peak turns out to be robust. Then the question arises which other features of the spectral functions are robust and not sensitive to the details of the spectral shape of the quasi-particle peak. In Ref. Bazavov et al. (2024) it was argued that the cumulants of the continuum-subtracted spectral function, defined as
@@
provide a robust characterization of the spectral function given the lattice data. The cumulants of the spectral function can be related to the cumulants of the correlation function for small $\tau$ Bazavov et al. (2024). In particular, $c_{2}$ is related to the second cumulant of the correlation function at $\tau=0$ Bazavov et al. (2024), which is the slope of the effective mass at small $\tau$. In the case of the Gaussian form of the dominant peak, $c_{2}=\Gamma_{G}$. In Tab. 4, Tab. 5 and Tab. 6, we show the values of $\sqrt{c_{2}}$ in lattice units. We see from the tables that, once the delta function at small $\omega$ is included in the fit, the values of $c_{2}$ obtained from the fit using the cut Lorentzian with $cut=2\Gamma_{0}$ and $cut=4\Gamma_{0}$, and using the Gaussian form agree. Thus, our results for $c_{2}$ can be used to constrain the parameters of other possible models for the spectral function and provide an alternative estimate of the thermal width.

### v2

@@ 3.1 Continuum-subtracted effective masses and in-medium bottomonium properties from simple ansätze
which for very large and very small $x$ behaves as $\exp(-2x/\delta)$. In the limit $\delta\rightarrow 0$, we recover the cut Lorentzian form. The parametrization of the spectral function given by Eq. 7, Eq. 8, and Eq. 10, which we will call it smooth-cut Lorentizan, can describe the spectral function obtained in the T-matrix approach Tang:2024dkz . This is shown in Fig. 8 for a 1P bottomonium state as an example. The parameter $\Gamma_{0}$ was obtained as the width at half maximum of the T-matrix spectral function, and the parameters $cut$ and $\delta$ were adjusted to obtain the best possible description of the T-matrix spectral functions. We also find that the Lorentzian form can describe the T-matrix spectral function very well in the interval $[M-{cut},M+{cut}]$, where ${cut}\simeq 4\Gamma_{0}$. A similar statement holds true for the spectral function of the static quark-antiquark pair obtained in hard thermal loop perturbation theory Burnier:2013fca .
@@
The in-medium mass shift $\Delta M(T)$ from wave-function-optimized operators is shown in Fig. 7. Interestingly, we see no mass shift within estimated errors for both the smooth-cut Lorentzian form and the cut Lorentzian form, for both lattice spacings. This is consistent with the previous findings in Refs. Larsen:2019bwy ; Larsen:2019zqv and also means that the absence of the mass shift is a robust result independent from the detailed form of the spectral function. We elaborate on this in Appendix B. The $\Delta M(T)$ for $\Upsilon$(1S) and $\chi_{b0}$(1P) states obtained from the correlators using Gaussian sources agree with the ones shown in Fig. 7 within errors.
@@
The in-medium width increases with the temperature and is larger for higher-lying states, i.e., it follows a hierarchy pattern corresponding to the hierarchy of bottomonium masses and sizes. In the case of the $\Upsilon$(1S), $\Upsilon$(2S), and $\chi_{b0}$(1P) states, the estimates of the thermal width obtained using the cut Lorentzian are significantly larger than the ones obtained using the smooth-cut Lorentzian. The difference is the largest for the $\Upsilon$(1S) state, where it is as large as a factor three. The difference in $\Gamma_{\rm med}(T)$ obtained from the two forms of the spectral function implies that the tail structure of the quasi-particle peak plays an important role in the observed decrease in the continuum-subtracted effective mass. The smaller the width, the larger the contribution from the tails of the dominant peak to the decrease in the effective masses. For this reason, the difference in $\Gamma_{\rm med}$ from using the two ansätze is the largest for the $\Upsilon$(1S). For the $\Upsilon$(3S) and the $\chi_{b0}$(2P), the difference in thermal width arising from the use of the two ansätze is much smaller, and for the $\chi_{b0}$(2P) state it is almost statistically insignificant. The reason for this is the following. When the width is large, the decrease in the effective masses is largely determined by the behavior of the spectral function around the peak, as explained in the Appendix B, and the tails of the quasi-particle peak do not contribute much to the observed decrease of the continuum-subtracted effective masses. The thermal widths obtained in this work are significantly smaller than the ones obtained in Refs. Larsen:2019bwy ; Larsen:2019zqv . The reason for this is the use of a simple Gaussian form for the quasi-particle peak in those studies, which is not physically motivated. For a Gaussian form of the peak, the slope of the continuum-subtracted effective masses is entirely determined by the width parameter. As discussed above, this is not the case for realistic spectral functions, for which the shape of the spectral function far away from the peak position can largely contribute to the slope of the continuum-subtracted effective masses.
@@ 4 Conclusions
In this paper we studied the in-medium properties of bottomonium using NRQCD with extended meson operators at two lattice spacings, $a=0.0602$ fm and $a=0.0493$ fm. We used temporal lattice extents $N_{\tau}=30-16$, covering the temperature range $T=133-251$ MeV. By studying the temperature dependence of bottomonium correlation functions, as well as through the comparison of the finite-temperature correlation function to the zero-temperature correlation function, we proposed a physically motivated parametrization of the in-medium spectral function that contains a dominant quasi-particle peak corresponding to the bottomonium states of interest. We found that all bottomonium states below the open-beauty threshold can exist as well-defined quasi-particle excitations up to $T=251$ MeV, though some of these excitations have large thermal width. For temperature $T\lesssim 167$ MeV the existence of bottomonium states and the values of their masses can be established without relying on any specific form of the spectral function. For the $\Upsilon$(1S) this is true even for $T\lesssim 182$ MeV. To obtain the in-medium bottomonium mass and width at higher temperatures, we considered several parameterizations of the dominant quasi-particle peak. We found that, irrespective of the details of these parameterizations, the in-medium bottomonium masses agree with the corresponding vacuum masses within errors. This is in contrast to the usual picture of bottomonium in quark-gluon plasma based on color screening, where there is significant shift in bottomonium masses and higher excited states are melted at temperatures somewhat above the crossover temperature. The remnants of confining interactions above the crossover temperature would be consistent with the existence of open charm hadrons suggested by the lattice QCD calculations of the generalized charm susceptibilities Mukherjee:2015mxc ; Bazavov:2023xzm . Furthermore, a recent lattice QCD study showed that topological objects underlying confinement exist well above the crossover temperature Mickley:2024vkm . In the future it would be interesting to extend these studies to higher temperatures and see the eventual melting of bottomonium states. This, however, should either involve the use of anisotropic lattices or relativistic formulations of the bottom quark since in present setup the condition $aM_{b}>0.9$ needed for NRQCD cannot be ensured for high temperatures.
@@
The estimates of the thermal width are quite sensitive to the detailed spectral shape of the quasi-particle peak. We considered two physically motivated choices of this spectral shape to obtain the thermal widths, which are shown in Fig. 9. These estimates of the thermal width are considerably smaller than the previous estimates from lattice NRQCD calculations that used a Gaussian form of the dominant peak Larsen:2019zqv ; Larsen:2019bwy , especially for the $\Upsilon$(1S). However, the two estimates of the thermal widths differ significantly, in particular for the $\Upsilon$(1S) state, where the differences can reach up to a factor of three. For the excited states, the differences in the thermal width obtained from the different spectral shapes are smaller, but, in general, estimating the thermal bottomonium width on the lattice better than a factor of two appears to be very challenging. Nevertheless, it would be very interesting to see how our estimates of the thermal width impact the phenomenological description of the nuclear suppression factor, $R_{AA}$, of the $\Upsilon$(1S), $\Upsilon$(2S), and $\Upsilon$(3S) states in heavy-ion collisions.
@@ Acknowledgements.
This work is supported partly by the National Key Research and Development Program of China under Contract No. 2022YFA1604900; the National Natural Science Foundation of China under Grants No. 12293064, No. 12293060 and No. 12325508, and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704, through Heavy Flavor Theory for QCD Matter (HEFTY) topical collaboration in Nuclear Theory, and within the framework of Scientific Discovery through Advance Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond. S. Meinel acknowledges support by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009913. W.-P. Huang acknowledges support from the Young Elite Scientists Sponsorship Program (Special Support for Doctoral Students) by CAST. R.L. was supported by the Ministry of Culture and Science of the State of Northrhine Westphalia (MKW NRW) under the funding code NW21-024-A (NRW-FAIR). This research used awards of computer time provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02- 05CH11231, on Frontier supercomputer in Oak Ridge Leadership Class Facility through ALCC Award, and the awards on JUWELS at GCS@FZJ, Germany and Marconi100 at CINECA, Italy. Computations for this work were carried out in part on facilities of the USQCD Collaboration, which are funded by the Office of Science of the U.S. Department of Energy. Some of the calculations were performed on the GPU cluster in the Nuclear Science Computing Center at CCNU (NSC3) and Wuhan supercomputing center.
@@ Appendix A Details of mass extractions in vacuum
In this Appendix, we describe the detailed procedure of fitting the vacuum correlators with an exponential ansätze to extract the corresponding mass parameters. We fitted the correlators to a single-exponential form within a range where the ground-state contribution dominates. Thus, careful selection of the fit interval is essential to eliminate contamination from higher states while maintaining statistical precision. To achieve this, we followed a procedure adapted from Refs. Fritzsch:2012wq ; Bahr2015Form . This procedure is exemplified in Fig. 10, using the $\chi_{b0}$(1P) case as an illustration. The upper end of the fit interval $\tau_{\text{max}}$ is determined based on the temporal separation where the signal-to-noise ratio of the effective mass $M_{\mathrm{eff}}(\tau)$ reaches a specified threshold, with the effective mass defined as $M_{\mathrm{eff}}(\tau)=-\partial_{\tau}\log C(\tau)$. In our analyses, we set a conservative criterion of $1\%$ for measurements employing Gaussian-smearing sources and $8\%$ for those using wave-function-optimized sources.
@@
To determine the lower end of the fit interval $\tau_{\text{min}}$, we initially performed a two-exponential fit to the correlators within a range where the two-state ansätze $f_{n}(\tau)=\sum_{i=0}^{n}A_{i}e^{-E_{i}\tau}$ with $n=1$ described the data well. This is illustrated by the solid colored lines in Fig. 10, with the two-state fits performed outside the gray-shaded area and ending at $\tau_{\text{max}}$. The fit parameters $A_{n=0,1}$ and $E_{n=0,1}$ were then used to reconstruct the excited-state contribution to the effective mass as $\delta M(\tau)=\log[f_{1}(\tau)/f_{1}(\tau+a)]-\log[f_{0}(\tau)/f_{0}(\tau+a)]$. We set $\tau_{\text{min}}$ to the smallest $\tau$ value where $\delta M(\tau)/E_{0}<25\%\times\delta_{M_{\mathrm{eff}}}(\tau)/M_{\mathrm{eff}}(\tau)$ is satisfied, with $\delta_{M_{\mathrm{eff}}}(\tau)$ representing the statistical uncertainty of $M_{\mathrm{eff}}$ at $\tau$. In this way, from $\tau_{\text{min}}$, the excited-state contamination (or, equivalently, the systematic uncertainty due to the fit ansätze) falls below the statistical uncertainty.
@@
For the $\chi_{b0}$(1P) case, $\tau_{\text{min}}$ is indicated by the dashed lines for wave-function optimized and Gaussian-smeared measurements respectively in Fig. 10, with the intervals $[\tau_{\text{min}},\tau_{\text{max}}]$ marked as the span of the colored bands. Within these intervals, the ground state dominates, allowing for a single-exponential fit to extract the information of the ground state. The error of the extracted mass is estimated using the Bootstrap method. The final central value of the mass was the median of the Bootstrap ensemble of fit results, with the error quoted as the 1-$\sigma$ width of this distribution. It is worth mentioning that for quantities such as mass differences, which depend on results from different fits, each fit was performed on the same Bootstrap sample. The final quantity for one sample is then computed using these fit results from the same sample, giving an approximate probability distribution. The mean value and the 1-$\sigma$ width of this distribution are then quoted.
Figure 10: The effective masses of $\chi_{b0}$(1P) from wave-function optimized (red) and Gaussian-smeared (blue) sources, enlarged from the results in Fig. 1. Bands (color-matched to the effective masses) indicate the masses extracted from single-exponential fits to the corresponding correlators. To determine the lower bound $\tau_{\text{min}}$ of the fit, initially, a two-state fit (colored curves) to the correlators is performed outside the gray-shaded area, terminating at $\tau_{\text{max}}$. The determined intervals $[\tau_{\text{min}},\tau_{\text{max}}]$ for the single-exponential fits are represented by the span of the colored bands.
@@ Appendix B Extractions of in-medium parameters at finite temperatures
As discusssed in Sec. 3 of the main text, the in-medium bottomonium parameters such as the widths and masses are encoded in the continuum-subtracted in-medium spectral function $\rho_{\rm med}(\omega,T)$. Furthermore, $\rho_{\rm med}(\omega,T)$ includes a quasi-particle peak containing information on the in-medium bottomonium masses and widths, and also a small contribution at frequencies well below the dominant peak controlling the behavior of the correlator at large $\tau$ around $1/T$. For the first part, we use the physically motivated cut Lorentzian ansätze or the smooth-cut Lorentzian ansätze, as discussed in the main text, while for the other part at $\omega$ well below the dominant-peak position, a single delta function is sufficient to describe the correlators at large $\tau$, shown as the first term of Eq. 11. In the main text it was argued that, for the cut Lorentzian, a reasonable choice for the cut is $4\Gamma_{0}$. Then an important question to address is how the obtained width depends on the choice of the parameter $cut$. A smaller value of $cut$ will result in a larger width. On the other hand the value of $cut$ cannot be too small because that would mean that the quasi-particle peak is not well defined, as other parts the spectral function provide a much larger contribution to the correlator. The value ${cut}=2\Gamma_{0}$ is probably the smallest value compatible with the existence of a quasi-particle peak, as the spectral function around the quasi-particle peak will decrease by a large factor before other structure in the spectral function becomes dominant. Therefore, we also considered the choice ${cut}=2\Gamma_{0}$ in our analysis as discussed below. For sufficiently low temperature, the thermal width can be too small to be resolved, given our statistical errors. In such cases, it makes sense to set the thermal width to zero, i.e., to assume a delta function for the quasi-particle peak. For sufficiently low temperature, the small contribution below the quasi-particle peak also does not play a role and can be neglected. This corresponds to fitting the effective masses by a constant. The results of these fits are shown in Tab. 2 and Tab. 3 for the $\Upsilon$ and $\chi_{b0}$ states respectively. In the cases of the $\Upsilon$(1S) state, the constant fit has worse $\chi^{2}/d.o.f$ compared to other fits only for $T\gtrsim 182$ MeV, while for other bottomonium states the constant fit has worse $\chi^{2}/d.o.f$ only for $T\gtrsim 163$ MeV.
@@
We also performed fits using the Gaussian form of the quasi-particle peak in the spectral function given by Eq. 6 to make contact with previous analyses Larsen:2019zqv ; Larsen:2019bwy , which are summarized in Tab. 6.
@@
The in-medium widths and mass shifts from different types of parametrizations of the in-medium spectral function $\rho_{\rm med}(\omega,T)$ are shown in Fig. 11 and Fig. 12. It can be seen from the comparison of open and filled points with the same shape in Fig. 11 that, for the excited bottomonium states, the widths extracted using the ansätze excluding a delta-function term tend to be systematically larger than those including the delta function, no matter what type of ansätze is used. This is especially obvious in the cases of high temperatures and for excited states. This indicates that the minor structure below the dominant peak of the spectral function matters for these cases, as expected. This delta peak is not well separated from the main peak structure of the spectral function, and interferences between them are possible. Thus, the tail of the main peak will compensate for the delta peak contribution if the delta term is excluded during the fits, leading to a larger width. In the case of the Gaussian fits the thermal width in this is defined as the width at the half maximum, which is related to the Gaussian width parameters as $\sqrt{2\ln 2}\Gamma_{G}$ Shi:2021qri . With this definition the thermal width from Gaussian fits width agrees with the cut Lorentzian results with $cut=2\Gamma_{0}$. As the cut increases to $4\Gamma_{0}$, the width extracted from cut Lorentzian ansätze decreases to about 50% $\sim$ 60% of the $cut=2\Gamma_{0}$ result, also stressing the importance of modeling the tail of the main peak structure of the spectral function. We also compare our estimates for the thermal width obtained from the cut Lorentzian form to estimates using the Gaussian ansätze for the dominant peak obtained in this analysis as well as in Ref. Larsen:2019bwy . Here again the width at half maximum was used to define the thermal width. Since the choice $cut=2\Gamma_{0}$ is the smallest value of cut still compatible with a well-defined quasi-particle peak and leads to the larger thermal width, the estimate of the width obtained from the Gaussian form of the dominant peak can be considered as an upper bound on the thermal bottomonium width. The in-medium mass shifts in Fig. 12 are all consistent with zero for all states under our consideration, irrelevant of the parametrizations of $\rho_{\rm med}(\omega,T)$.
@@
The detailed choice of the quasi-particle peak has significant effect on the estimated thermal bottomonium width, while the position of the peak turns out to be robust. Then the question arises which other features of the spectral functions are robust and not sensitive to the details of the spectral shape of the quasi-particle peak. In Ref. Bazavov:2023dci it was argued that the cumulants of the continuum-subtracted spectral function, defined as
@@
provide a robust characterization of the spectral function given the lattice data. The cumulants of the spectral function can be related to the cumulants of the correlation function for small $\tau$ Bazavov:2023dci . In particular, $c_{2}$ is related to the second cumulant of the correlation function at $\tau=0$ Bazavov:2023dci , which is the slope of the effective mass at small $\tau$. In the case of the Gaussian form of the dominant peak, $c_{2}=\Gamma_{G}$. In Tab. 4, Tab. 5 and Tab. 6, we show the values of $\sqrt{c_{2}}$ in lattice units. We see from the tables that, once the delta function at small $\omega$ is included in the fit, the values of $c_{2}$ obtained from the fit using the cut Lorentzian with $cut=2\Gamma_{0}$ and $cut=4\Gamma_{0}$, and using the Gaussian form agree. Thus, our results for $c_{2}$ can be used to constrain the parameters of other possible models for the spectral function and provide an alternative estimate of the thermal width.